A boat floating in fresh water displaces water weighing 32.9 kN. (a) What is the weight of the water this boat displaces when floating in salt water of density 1.14 × 103 kg/m3? (b) What is the absolute value of the difference between the volume of fresh water displaced and the volume of salt water displaced?
Answer #1

(a). The weight of salt water displaced is still 32.9kN.
(b). The difference in volume displaced:
(32900N/9.8) = 3357.14 kg displacement.
In fresh water, this would be 3357.14 litres.
In salt water, this would be (3357.14/1.14) = 2944.86 litres.
so absolute volume difference = 3357.14 - 2944.86 = 412.278 litres.